OpenStudy (anonymous):

Let A be a 2 x 2 such that [2] is an eigenvector for [3] A with eigenvalue 1, and [1] is an eigenvector for A [2] with eigenvalue 1/2. If v = [-1] [1 ] Compute A^4 V? Help please.

OpenStudy (jlg030597):

Is this calculus

OpenStudy (loser66):

linear algebra

OpenStudy (jlg030597):

Oh so Algebra 2

OpenStudy (jlg030597):

I sctuslly want to learn how this works im in Geometry so learning ahead can't be that bad

OpenStudy (kc_kennylau):

(I just did some research myself from wikipedia so i may not be reliable) Maybe let the matrix be \(\left(\begin{matrix}a&b\\c&d\end{matrix}\right)\)?

OpenStudy (jlg030597):

Whats a matrix

OpenStudy (kc_kennylau):

I thought you learnt what's a matrix... Well before you learn eigenvalues and eigenvectors you must learn matrix right?

OpenStudy (anonymous):

so i have to find A using the eigenvectors and eigenvalues?

OpenStudy (jlg030597):

Im only 10th grade

OpenStudy (kc_kennylau):

but I thought you are learning ahead? P.S. no bragging by I knew matrix since i was in 5th grade or so

OpenStudy (kc_kennylau):

but*

OpenStudy (loser66):

hope that helps

OpenStudy (jlg030597):

@kc_kennylau Im happy for you but Im in tenth grade know what vectors are

OpenStudy (kc_kennylau):

Eigenvector is not exactly related to vector :)

OpenStudy (anonymous):

How do i know i should find [P^-1] and D^4?

OpenStudy (jlg030597):

I was bragging that i knew what vectors are lol im studying on Khan Academy

OpenStudy (kc_kennylau):

wait @jlg030597 I thought you were the question owner lol fml

OpenStudy (kc_kennylau):

@jan4712 but there ain't P and D...

OpenStudy (jlg030597):

Nope Jan is I wanted to learn how it worked

OpenStudy (loser66):

@jan4712 if you didn't know, now, you know. That's the process to find out exponential matrix

OpenStudy (anonymous):

sorry i was referring to loser66. so D^n for A^n?

OpenStudy (loser66):

\[A^n \neq D^n\] \[A^n = P^- * D^n*P\] 3 matrices time together, and the order MUST be that.

OpenStudy (jlg030597):

Oh my Im so confused now

OpenStudy (loser66):

@jlg030597 you have to have some basis to solve this problem

OpenStudy (anonymous):

i get it. thank you!

OpenStudy (loser66):

np